OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 1 centimeters/minute while the area of the triangle is increasing at a rate of 5 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 10 centimeters and the area is 88 square centimeters?

OpenStudy (bahrom7893):

why don't you try this one now.

OpenStudy (bahrom7893):

i solved like 5 of these.

OpenStudy (anonymous):

Ok, I'll try it and post the answer that I get and you let me know if I got it right.

OpenStudy (bahrom7893):

ok

OpenStudy (anonymous):

-.76?

OpenStudy (bahrom7893):

huh? It can't be negative, the area is increasing.

OpenStudy (bahrom7893):

A = (1/2)b*h dA/dt = (1/2)(b*dh/dt + h*db/dt)

OpenStudy (bahrom7893):

dA/dt = 5 dh/dt = 1 h=10 A=88 db/dt - ?

OpenStudy (bahrom7893):

10 = b*1 + 10*db/dt

OpenStudy (bahrom7893):

10 = b + 10(db/dt)

OpenStudy (bahrom7893):

Now all we gotta do is figure out b when A = 88 and h = 10 A = (1/2)b*h 88 = (1/2)b*10 88 = 5b b = 88/5

OpenStudy (bahrom7893):

10 = b + 10(db/dt) 10 = 88/5 + 10(db/dt) 50 = 88 + 50(db/dt) 50-88 = 50(db/dt) -38 = 50(db/dt) db/dt = -0.76. You were right!

OpenStudy (bahrom7893):

hmm I guess the height is increasing pretty fast, so that the area manages to increase while the base decreases.

OpenStudy (bahrom7893):

Good job!

OpenStudy (anonymous):

Yay! Thanks!!!